OpenStudy (babyx3boo):

a^-2+a^-1b-b^2/ a^2b^-1 anyone knows how to do this? :o

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i think i do i take it those are negative exponents right?

OpenStudy (anonymous):

You're just simplifying, correct?

OpenStudy (misty1212):

\[\huge \frac{a^{-2}+a^{-1}b-b^2}{a^2b^{-1}}\]??

OpenStudy (misty1212):

is that it or not quite?

OpenStudy (anonymous):

Parenthesis would be really helpful here\[(a ^{-2}+a^{-1})(b-b^{-2})\div (a^{2})(b^{-1})\]

OpenStudy (anonymous):

If that's what you're going for

OpenStudy (babyx3boo):

@misty1212 yes that's right & @Occisor i'm guessing it just says perform the operation.

OpenStudy (jhannybean):

Divide each term in the numerator by the denominator.

OpenStudy (jhannybean):

\[\large \frac{a^{-2}}{a^2b^{-1}}+\frac{a^{-1}b}{a^2b^{-1}} -\frac{b^2}{a^2b^{-1}}\]

OpenStudy (misty1212):

you get nothing real nice out of this, just three terms as @Jhannybean wrote

OpenStudy (jhannybean):

Remember this rule? : \(\large \frac{x^m}{x^n} = x^{m-n}\) Take for example \(x=a\). Now \(a\) is being divided by itself but with 2 different powers, \(n=-2~,~ m=2\). So what you would do here is.... \[\large \frac{a^{-2}}{a^2} = a^{-2-2} = a^{-4} = \frac{1}{a^4}\]

OpenStudy (jhannybean):

@babyx3boo is anything making sense?...

OpenStudy (babyx3boo):

sorta; it's just weird doing what you said i got 1/a^-4b^-1 + 1/a^-3b^2-b^3/a^2

OpenStudy (bee_see):

You have to change the sign if you move the variable from the numerator to denominator or denominator to numerator

OpenStudy (bee_see):